# CoVoMix: Advancing Zero-Shot Speech Generation for Human-like Multi-talker Conversations

### v1

@@ Abstract
Recent advancements in zero-shot text-to-speech (TTS) modeling have led to significant strides in generating high-fidelity and diverse speech. However, dialogue generation, along with achieving human-like naturalness in speech, continues to be a challenge in the field. In this paper, we introduce CoVoMix: Conversational Voice Mixture Generation, a novel model for zero-shot, human-like, multi-speaker, multi-round dialogue speech generation. CoVoMix is capable of first converting dialogue text into multiple streams of discrete tokens, with each token stream representing semantic information for individual talkers. These token streams are then fed into a flow-matching based acoustic model to generate mixed mel-spectrograms. Finally, the speech waveforms are produced using a HiFi-GAN model. Furthermore, we devise a comprehensive set of metrics for measuring the effectiveness of dialogue modeling and generation. Our experimental results show that CoVoMix can generate dialogues that are not only human-like in their naturalness and coherence but also involve multiple talkers engaging in multiple rounds of conversation. These dialogues, generated within a single channel, are characterized by seamless speech transitions, including overlapping speech, and appropriate paralinguistic behaviors such as laughter. Audio samples are available at https://aka.ms/covomix.
@@ 1 Introduction
Zero-shot Text-to-speech (TTS) technology aims to create human-like natural speech with voice characteristics prompted by the context. Recent deep learning advancements have significantly improved synthesized speech quality, especially in formal reading scenarios [1; 2; 3]. However, TTS systems still struggle with rendering spontaneous-style speech and managing seamless transitions during conversations—common occurrences in everyday human communication [4; 5]. On the one hand, spontaneous-style speech encompasses phenomena like filled pauses, interjections, repairs, repetitions, and laughter, which lend human-realistic naturalness to spoken language [6]. On the other hand, in natural conversations, speakers intuitively time their speech, determining when to speak and when to yield, resulting in seamless transitions with appropriate overlaps or moments of silence [7; 8]. Considering these conversational features, we summarize three main challenges in generating spontaneous dialogues.
First, the scarcity of high-quality, spontaneous conversational datasets, along with the difficulty in segmenting paralinguistic behaviors, continues to be a significant obstacle in the field. Spontaneous behavior and non-verbal expressions such as laughter, receive insufficient attention in speech synthesis. Existing high-quality datasets for spontaneous and conversational speech are relatively small and involve a limited number of speakers [9]. Identifying and segmenting these paralinguistic features, as highlighted by studies [10; 11], poses difficulties. Models that require pre-alignment necessitate sophisticated manual annotation. Otherwise, the low-quality data can adversely impact performance, particularly in TTS tasks [12].
@@
    To the best of our knowledge, it is the first attempt at zero-shot, human-like multi-speaker multi-round dialogue generation model. We propose 1) a simultaneous multi-stream semantic token prediction from a dialogue text; and 2) a multi-talker flow-matching based acoustic model for generating a mixed mono mel-spectrogram given multiple contexts. It is capable of generating single-channel multi-round dialogue containing multiple speakers concurrently, enabling simultaneous timbre cloning of multiple speakers in zero-shot scenarios.
@@
    The model, trained on conversational speech, transforms continuous speech features into discrete tokens, capturing paralinguistic behaviors. An autoregressive model is utilized to convert dialogue text, delimited by speaker tags, to these discrete token sequences for each speaker. The overlapping segments, specifically the duration of overlap among speakers, are learned from the training data. Consequently, during the inference stage, these overlaps are automatically generated without the need for any special symbols in the input text.
@@
  4. 4.
    We employ the Fisher dataset for this study, which was curated for robust speech recognition. Our approach includes a comprehensive strategy for processing this dataset, including both training and evaluation for monologue and dialogue speech. The data processing script, along with the model training codes and evaluation tools, will be publicly available.
@@ 2 Related Work
2.1 Text-to-Speech as Language Modeling
With the advancement in neural codecs for speech, many recent studies have explored token-based language modeling for text-to-speech task. VALL-E [2], a zero-shot text-to-speech model, is a text conditioned language model trained on EnCodec tokens [17]. SPEAR-TTS [18] casts TTS as two tasks: text to high-level semantic tokens and semantic tokens to low-level acoustic tokens. Base-TTS [19] is the largest TTS model to date, which is trained on 100K hours data and utilizes discrete codes extracted from a WavLM Self-Supervised Learning (SSL) speech model [20] as speech representation. HuBERT [21], another model within the SSL framework, serves as the foundational architecture for the Generative Spoken Language Modeling (GSLM) family [22; 8]. The GSLM models leverage HuBERT units specifically for the task of speech continuation, distinguishing themselves by operating directly on raw audio signals without the need for any labels or accompanying text.
2.2 Flow Matching
Continuous Normalizing Flows (CNF) [23] is a deterministic framework that is capable of encompassing the probability paths modeled by diffusion processes [24]. It learns the transformation from normal distribution $p_{0}$ to data distribution $p_{1}$, and it parameterizes a time-dependent vector field $v_{t}$ that pushes samples from prior towards the target distribution [25; 26].
Flow Matching (FM) [25] is a simulation-free approach for training CNFs at scale based on regressing vector fields of fixed conditional probability paths. The relationship between the vector field and the flow $\phi$ is defined via an ordinary differential equation (ODE) $d\phi_{t}(y)=v_{t}(\phi_{t}(y))dt,\phi_{0}(y)=y$ [25]. Benefiting from flow matching models, text-to-speech models, such as VoiceFlow [27], MatchaTTS [28] and Voicebox  [26], can generate high-quality speech efficiently.
@@ 2.3 Dialogue Generation
With the application of large-scale data and language modeling, generating continuous conversations becomes possible. dGSLM [8] is built on HuBERT semantic token sequence for speech continuation task, which generates two-channel spoken dialogue auto-regressively. However, the textless design prevents dGSLM from controlling the content of the generated speech, resulting in output that is often less intelligible.
SoundStorm [29] is an iterative generative method that converts semantic tokens into acoustic audio tokens. It can perform monologue and dialogue synthesis. However, it synthesizes dialogue sequentially and in reading style, lacking any spontaneous behaviors or instances of overlapping speech.
CHATS [30], similar to dGSLM [8], is an auto-regressive dialogue generation model, which transforms written dialogues into spoken conversations. It is conditioned on the phonetic transcription and can generate meaningful utterances on the speaker side. However, it is unable to control the generation of spontaneous behaviors for both speakers. Its language model is limited to processing a fixed number of dialogue rounds and can only accommodate a predetermined number of speakers. This restriction narrows its potential use cases.
3 Method
3.1 Problem Formulation
Zero-shot speech generation is a task where a model synthesizes speech in a target voice that was not present in its training data. This task requires only a transcript of what is to be spoken and a speech prompt—a brief sample recording of the target voice. It is generally achieved by in-context learning with a dataset of transcribed speech $\{x,y\}$ where ${y}$ and ${x}$ denote speech utterances and their transcriptions, respectively. Zero-shot multi-talker conversational speech synthesis is designed to generate the voices of multiple speakers simultaneously, based on their transcriptions and prompts. Our approach diverges from the traditional method of synthesizing each voice individually and then concatenating them. The goal is to capture the dynamic nature of real conversations, where participants may speak over each other or respond spontaneously with interjections such as laughter.
Our proposed method, CoVoMix, consists of a multi-stream text-to-semantic model, an acoustic model and a vocoder. The text-to-semantic model first generates multi-stream semantic token sequences for each speaker, given the dialogue transcription. Then the acoustic model transforms these semantic sequences into a mixed mel-spectrogram. A vanilla HiFi-GAN vocoder [31] finally synthesizes mono-channel multi-round dialogue from the mel-spectrogram. We utilize a conversational dataset $D=\{x,y\}$ for training, where $y=[y^{1},y^{2}]$ represents a stereo dialogue featuring two speakers, and $x$ corresponds to the text transcription annotated with speaker tags.
@@ 3.2 Multi-stream Text-to-Semantic Model
The multi-stream text-to-semantic model is a sequence-to-sequence model based on encoder-decoder architecture. It takes in a text token sequence generated by a BERT text tokenizer  [32], augmented with special tokens denoting speaker transitions and interjections. The output comprises a multi-stream semantic token sequence. For this study, we focus on a dual-stream setup for a dialogue between two speakers. We employ a pre-trained HuBERT model  [^1] https://github.com/facebookresearch/fairseq/tree/main/examples/textless_nlp/dgslm/hubert_fisher [8] as a speech tokenizer to extract the clustered discrete HuBERT hidden units as semantic token sequences and process two channels of waveform, separately. If the dialogues are captured in a single-channel recording, it is necessary to perform speaker separation to produce a dual-channel waveform in our approach. The process of semantic token extraction operates at the frame level, with a time shift of 20 milliseconds, resulting in the presence of duplicated tokens within the sequence. We train this model on a paired speech-text dataset with cross-entropy loss, as
@@ 3.3 Acoustic Model
At each timestamp $t\in[0,1]$, a lookup table first embeds the semantic token sequence $s=[s^{1},s^{2}]$ into $s_{emb}=[s_{emb}^{1},s_{emb}^{2}]$ for two speakers. We extract the corresponding mixed mel-spectrogram $m$ and individual mel-spectrogram $[m^{1},m^{2}]$ for each speaker of dialogue $y$. We randomly choose a mask. The masked part $\tilde{m}=m\odot mask$ is to predict, while the seen part $m_{ctx}=[m^{1}\odot(1-mask),m^{2}\odot(1-mask)]$ is considered as prompt.
At each flow step $t$, we sample $w=(1-(1-\sigma_{min})t)\tilde{m_{0}}+tm$)], where $\sigma_{min}$ is a hyper-parameter controlling deviation, $\tilde{m_{0}}$ is sampled from $\mathcal{N}(m|0,\mathbf{I})$. Then, the sample $w$ at flow step $t$, the acoustic prompt $m_{ctx}$, and semantic embedding sequences $s_{emb}$ are concatenated frame-by-frame to obtain an input matrix $W_{input}$. Conditional Flow Matching (CFM) is a per-example training objective, which provides equivalent gradients and does not require explicit knowledge of the intractable target vector field. Therefore, we train the acoustic model to learn the mixed mel-spectrogram with objective as in Eq.2, with $v_{t}(w,m_{ctx},s_{emb};\theta)$ the transformer output with flow $w$ at step $t$.
@@
We also use classifier-free guidance during training for flow-matching model, which is to trade off mode coverage and sample fidelity [26; 33]. During training, the acoustic prompt $m_{ctx}$ and semantic sequences $s_{emb}$ are dropped with $p_{uncond}$. During inference, we use the modified vector field $\tilde{v}_{t}(w,m_{ctx},s_{emb};\theta)$ written in Equation 3 to replace $v_{t}(w,m_{ctx},s_{emb};\theta)$, where $\alpha$ is a hyperparameter controlling the strength of guidance.
@@ 4.1 Data Preparation
The dataset used in this work is Fisher Dataset [34], which is a telephone conversation dataset with 2,000h English conversations about various topics. Each dialogue was recorded in two channels with an 8kHz sample rate and an average duration of 10 minutes. We randomly divide the Fisher dataset into train/valid/test sets with 97/1/2/ split. Each set has different speakers.
The data preparation is different for monologue and dialogue. For monologue, following Nemo [35] script [^2] https://gitlab.nrp-nautilus.io/ar-noc/nemo/-/blob/master/scripts/process_fisher_data.py, we slice long dialogues into smaller mono-channel samples, and concatenate them to match minimum duration. We also prepare the corresponding transcripts, extract the corresponding mel spectrogram and semantic token sequence for each sample. For dialogue, we slice long dialogues into shorter, stereo-channel dialogues containing at least two utterances from distinct speakers. We ensure that the first and last sentences of each processed dialogue do not overlap with other dialogues, thus avoiding any extraneous content in the transcriptions. We organize the multi-round dialogue transcript chronologically by the start time of each utterance, delimited by the ‘[spkchange]’ token. The HuBERT speech tokenizer is employed to extract the semantic tokens for each channel. Additionally, we extract the mel-spectrogram from the mixed waveform. The detailed algorithm for dialogue data preparation is described in Appendix C.
4.2 Model Configarations

### v2

@@ Abstract
Recent advancements in zero-shot text-to-speech (TTS) modeling have led to significant strides in generating high-fidelity and diverse speech. However, dialogue generation, along with achieving human-like naturalness in speech, continues to be a challenge. In this paper, we introduce CoVoMix: Conversational Voice Mixture Generation, a novel model for zero-shot, human-like, multi-speaker, multi-round dialogue speech generation. CoVoMix first converts dialogue text into multiple streams of discrete tokens, with each token stream representing semantic information for individual talkers. These token streams are then fed into a flow-matching based acoustic model to generate mixed mel-spectrograms. Finally, the speech waveforms are produced using a HiFi-GAN model. Furthermore, we devise a comprehensive set of metrics for measuring the effectiveness of dialogue modeling and generation. Our experimental results show that CoVoMix can generate dialogues that are not only human-like in their naturalness and coherence but also involve multiple talkers engaging in multiple rounds of conversation. This is exemplified by instances generated in a single channel where one speaker’s utterance is seamlessly mixed with another’s interjections or laughter, indicating the latter’s role as an attentive listener. Audio samples are available at https://aka.ms/covomix.
@@ 1 Introduction
Zero-shot Text-to-speech (TTS) technology aims to create human-like natural speech with voice characteristics prompted by the context. Recent deep learning advancements have significantly improved synthesized speech quality, especially in formal reading scenarios [1, 2, 3]. However, TTS systems still struggle with rendering spontaneous-style speech and managing seamless transitions during conversations—common occurrences in everyday human communication [4, 5]. On the one hand, spontaneous-style speech encompasses phenomena like filled pauses, interjections, repairs, repetitions, and laughter, which lend human-realistic naturalness to spoken language [6]. On the other hand, in natural conversations, speakers intuitively time their speech, determining when to speak and when to yield, resulting in seamless transitions with appropriate overlaps or moments of silence [7, 8]. Overlapping speech, defined as more than one person is speaking, can easily exceed 20%, in informal gathering conversational speech [9]. Considering these conversational features, we summarize three main challenges in generating spontaneous dialogues.
First, the scarcity of high-quality, spontaneous conversational datasets, along with the difficulty in segmenting paralinguistic behaviors, continues to be a significant obstacle in the field. Spontaneous behavior and non-verbal expressions such as laughter, receive insufficient attention in speech synthesis. Existing high-quality datasets for spontaneous and conversational speech are relatively small and involve a limited number of speakers [10]. Identifying and segmenting these paralinguistic features, as highlighted by studies [11, 12], poses difficulties. Models that require pre-alignment necessitate sophisticated manual annotation. Otherwise, the low-quality data can adversely impact performance, particularly in TTS tasks [13].
@@
    To the best of our knowledge, it is the first attempt at zero-shot, human-like, multi-talker conversational mixed speech generation. We propose 1) a simultaneous multi-stream semantic token prediction, with each stream representing an individual talker, from dialogue text; and 2) a multi-talker flow-matching based acoustic model for generating a mixed mono mel-spectrogram given multiple contexts. It is capable of generating single-channel multi-round dialogue containing multiple speakers concurrently, enabling simultaneous timbre cloning of multiple speakers in zero-shot scenarios.
@@
    We employ the Fisher dataset [18] for this study, which was curated for robust speech recognition. Our approach includes a comprehensive strategy for processing this dataset, including both training and evaluation for monologue and dialogue speech. The data processing script, along with the model training codes and evaluation tools, will be publicly available.
@@ 2 Related Work
2.1 Zero-shot Text-to-Speech
The goal of Zero-shot TTS is to synthesize speech in a target voice which was unseen during training, given only the target transcript and a short reference of the target voice  [19, 20, 1, 21] . Zero-shot TTS systems are generally divided into two categories: (i) Diffusion-based Zero-Shot TTS [3, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32] and (ii) Neural Codec-based Zero-shot TTS [2, 33, 34, 35, 36].
Diffusion-based Zero-shot TTS models handle the problem in a non-auto regressive manner and have shown excellent performance in audio generation tasks [25, 37, 38]. Many previous works, such as [22, 29, 39], use log Mel spectrograms as intermediate features and generate speech waveforms using high-quality vocoders. For instance, DiffVoice [40] employs a VAE-GAN autoencoder [41] to encode the Mel-Spectrogram into a latent space, jointly modeling phoneme duration and Mel-spectrogram. FastSpeech [42] generates mel-spectrograms in parallel for faster inference, managing alignment between phoneme sequence and generated spectrogram with an explicit length regulator and duration predictor model. Flow matching training is a method that is closely related to Diffusion models offering simpler trajectories and requiring fewer function evaluations during inference [43, 44, 31]. Flow Matching (FM) [43] is a simulation-free approach for training continuous normalizing flows (CNFs) at scale based on regressing vector fields of fixed conditional probability paths. The relationship between the vector field and the flow $\phi$ is defined via an ordinary differential equation (ODE) $d\phi_{t}(y)=v_{t}(\phi_{t}(y))dt,\phi_{0}(y)=y$ [43]. Benefiting from flow matching models, text-to-speech models, such as VoiceFlow [45], MatchaTTS [38] and Voicebox  [46], can generate high-quality speech efficiently.
On the other side, Neural Codec-based methods formulate the TTS problem as a token-based language modeling task [2, 47]. VALL-E [2], a zero-shot text-to-speech model, is a text conditioned language model trained on EnCodec tokens [47]. SPEAR-TTS [48] is similar to AudioLM [49] which carries out the speech generation process in two steps: first, it maps the text into discrete semantic tokens, then, in the second step, it converts the semantic tokens into acoustic tokens. In the recently proposed BASE-TTS [50], the authors propose to model the joint distribution of text tokens and discrete speech representations referred to as speechcodecs followed by a convolution-based decoder which converts these speechcodes into waveforms in an incremental, streamable manner. NaturalSpeech3 [51] combines codecs and diffusion modeling, achieving significant improvements in speech quality, prosody, intelligibility, and scalability with a 1B-parameter model trained on 200K hours of data. It decomposes speech waveforms into content, prosody, timbre, and acoustic details, reconstructing speech from these disentangled representations using a factorized diffusion model.
@@ 2.2 Dialogue Generation
dGSLM [8] represents the pioneering textless model for generating naturalistic spoken dialogues. It utilizes a dual-tower transformer with cross-attention as its architectural backbone and leverages HuBERT semantic token sequence as its input for the speech continuation task. This model generates two-channel spoken dialogue auto-regressively, without reliance on text or labels. However, its textless nature constrains its ability to direct the content of the speech it produces, occasionally leading to less intelligible outputs.
CHATS [52], while based on the same architectural principles as dGSLM, is designed to convert written dialogues into spoken conversations. It is capable of generating speech for both speaker and listener sides, conditioning on speaker ID, phoneme sequence, and context from the speaker’s side, without requiring transcriptions for spontaneous behaviors or laughter. However, it does not support the capabilities of zero-shot voice cloning, relying solely on speaker ID for retaining speaker characteristics.
SoundStorm [33], on the other side, is an iterative generative method that converts semantic tokens into acoustic audio tokens. It can perform zero-shot monologue and dialogue synthesis. Yet, the synthesized dialogue is generated in a sequential manner and thus sounds less realistic, lacking any spontaneous behaviors or instances of overlapping speech.
3 CoVoMix
Zero-shot speech generation is a task where a model synthesizes speech in a target voice that was not present in its training data. This task requires only a transcript of what is to be spoken and a speech prompt—a brief sample recording of the target voice. It is generally achieved by in-context learning with a dataset of transcribed speech $\{x,y\}$ where ${y}$ and ${x}$ denote speech utterances and their transcriptions, respectively. Zero-shot multi-talker conversational speech synthesis is designed to generate the voices of multiple speakers simultaneously, based on their transcriptions and prompts. Our approach differs from the traditional method in which each voice is synthesized individually, and then concatenated to form a dialogue. Our goal in this work is to capture the dynamic nature of real conversations, where participants may speak over each other or respond spontaneously with interjections such as laughter.
Our proposed CoVoMix, shown in Figure 1, consists of a multi-stream text-to-semantic model, an acoustic model and a vocoder. The text-to-semantic model first generates multi-stream semantic token sequences for each speaker, given the dialogue transcription. Then the acoustic model transforms these semantic sequences into a mixed mel-spectrogram. A vanilla HiFi-GAN vocoder [53] finally synthesizes mono-channel multi-round dialogue from the mel-spectrogram. We utilize a conversational dataset $D=\{x,y\}$ for training, where $y=[y^{1},y^{2}]$ represents a stereo dialogue featuring two speakers, and $x$ corresponds to the text transcription annotated with speaker tags.
@@ 3.1 Multi-stream Text-to-Semantic Model
The multi-stream text-to-semantic model is a sequence-to-sequence model based on encoder-decoder architecture. It takes in a text token sequence generated by a BERT text tokenizer  [54], augmented with special tokens denoting speaker transitions and interjections. The output comprises a multi-stream semantic token sequence. For this study, we focus on a dual-stream setup for a dialogue between two speakers. We employ a pre-trained HuBERT model  [^1] https://github.com/facebookresearch/fairseq/tree/main/examples/textless_nlp/dgslm/hubert_fisher (MIT License) [8] as a speech tokenizer to extract the clustered discrete HuBERT hidden units as semantic token sequences and process two channels of waveform, separately. If the dialogues are captured in a single-channel recording, it is necessary to perform speaker separation to produce a dual-channel waveform in our approach. The process of semantic token extraction operates at the frame level, with a time shift of 20 milliseconds, resulting in the presence of duplicated tokens within the sequence. We train this model on a paired speech-text dataset with cross-entropy loss, as
@@ 3.2 Acoustic Model
At each timestamp $t\in[0,1]$, a lookup table first embeds the semantic token sequence $s=[s^{1},s^{2}]$ into $s_{emb}=[s_{emb}^{1},s_{emb}^{2}]$ for two speakers. We extract the corresponding mixed mel-spectrogram $m$ and individual mel-spectrogram $[m^{1},m^{2}]$ for each speaker of dialogue $y$. We randomly choose a mask. The masked part $\tilde{m}=m\odot mask$ is to be predicted, while the seen part $m_{ctx}=[m^{1}\odot(1-mask),m^{2}\odot(1-mask)]$ is considered as prompt.
At each flow step $t$, we sample $w=(1-(1-\sigma_{min})t)\tilde{m_{0}}+tm$)], where $\sigma_{min}$ is a hyper-parameter to control deviation and $\tilde{m_{0}}$ is sampled from $\mathcal{N}(m|0,\mathbf{I})$. Then, the sample $w$ at flow step $t$, the acoustic prompt $m_{ctx}$, and semantic embedding sequences $s_{emb}$ are concatenated frame-by-frame to obtain an input matrix $W_{input}$. Conditional Flow Matching (CFM) [43] is a per-example training objective, which provides equivalent gradients and does not require explicit knowledge of the intractable target vector field. Therefore, we train the acoustic model to learn the mixed mel-spectrogram with objective as in Eq.2, where $v_{t}(w,m_{ctx},s_{emb};\theta)$ is the transformer output with flow $w$ at step $t$.
@@
We also use classifier-free guidance, a method to trade off mode coverage and sample fidelity [46, 55], in the training for flow-matching model. During training, the acoustic prompt $m_{ctx}$ and semantic sequences $s_{emb}$ are dropped with $p_{uncond}$. During inference, we use the modified vector field $\tilde{v}_{t}(w,m_{ctx},s_{emb};\theta)$ shown in Equation 3 to replace $v_{t}(w,m_{ctx},s_{emb};\theta)$, where $\alpha$ is a hyperparameter controlling the strength of guidance.
@@ 4.1 Data Preparation
The dataset used in this work is Fisher dataset [18], which is a telephone conversation dataset with 2,000h English conversations about various topics. Each dialogue was recorded in two channels with an 8kHz sample rate and an average duration of 10 minutes. We randomly divide the Fisher dataset into train/valid/test sets with 97/1/2 split. Each set has different speakers.
The data preparation is different for monologue and dialogue. For monologue, following Nemo [56] script, [^2] https://gitlab.nrp-nautilus.io/ar-noc/nemo/-/blob/master/scripts/process_fisher_data.py (Apache License 2.0) we slice long dialogues into smaller mono-channel samples and concatenate them to meet the minimum duration requirement, which is set to 10 seconds by default. We prepare the corresponding transcripts, extract the mel-spectrogram and semantic token sequence for each sample. Spontaneous behavior such as laughter is labeled by [laughter] token in the transcription. For dialogue, we slice long dialogues into shorter, stereo-channel dialogues containing at least two utterances from distinct speakers. We ensure that the first and last sentences of each processed dialogue do not overlap with other dialogues, thus avoiding any extraneous content in the transcriptions. Motivated by serialized output training in speech recognition task [57, 58], we organize the multi-round dialogue transcript chronologically by the start time of each utterance. Two neighboring same-speaker utterances are concatenated directly, while different speakers’ utterances are separated by [spkchange] token, without explicit processing for overlap labelling. A dialogue transcription preparation example is shown in Figure 2. The HuBERT speech tokenizer is employed to extract the semantic tokens for each channel. Additionally, we mix audio with two channels and extract the mel-spectrogram from the mixed waveform. The detailed algorithm for dialogue data preparation is described in Appendix E.
Figure 2: Dialogue transcription preparation. To better demonstrate our method, we use | and emoji to represent [spkchange] and [laughter] tokens.
4.2 Model Configurations